the sum of the ages of two sisters is 21.
the sum of the squares of their ages is 233.
what are the two ages?
Their ages can be represented by x and y.
x+y=21
x^2+y^2=233

Answer by mananth(16946)

x+y=21
x^2+y^2=233
x^2+y^2+2xy = 233+2xy
(x+y)^2 = 233+2xy
21^2= 233 +2xy
441-233= 2xy
208=2xy
104 = xy
y= 104/x
x+104/x=21
x^2+104 / x = 21
x^2+104=21x
x^2-21x+104=0
x^2-8x-13x+104=0
x(x-8)-13(x-8)=0
(x-8)(x-13)=0
x=8 Or x=13
the ages are 8 and 13